# Improving Model Factuality with Fine-grained Critique-based Evaluator

### v1

@@ Abstract
Factuality evaluation aims to detect factual errors produced by language models (LMs) and hence guide the development of more factual models. Towards this goal, we train a factuality evaluator, FenCE, that provides LM generators with claim-level factuality feedback. We conduct data augmentation on a combination of public judgment datasets to train FenCE to (1) generate textual critiques along with scores and (2) make claim-level judgment based on diverse source documents obtained by various tools. We then present a framework that leverages FenCE to improve the factuality of LM generators by constructing training data. Specifically, we generate a set of candidate responses, leverage FenCE to revise and score each response without introducing lesser-known facts, and train the generator by preferring highly scored revised responses. Experiments show that our data augmentation methods improve the evaluator’s accuracy by 2.9% on LLM-AggreFact. With FenCE, we improve Llama3-8B-chat’s factuality rate by 14.45% on FActScore, outperforming state-of-the-art factuality finetuning methods by 6.96%.
@@ 1 Introduction
To avoid latency issues in inference, prior studies try to train the generator to output more factual responses, by (1) preferring generation candidates with higher factuality scores (Tian et al., 2024), which is limited by the generator’s capabilities, or (2) preferring responses with false information corrected (Kang et al., 2024a), which is prone to introducing lesser-known facts that are not well memorized in pretraining. As shown by recent work (Gekhman et al., 2024; Ghosal et al., 2024), such preference training reinforces the model to generate information that are not well memorized, and hence could even hurt factuality. Furthermore, the methods either leverage proprietary models that have restricted terms of use in evaluation, or prompt the generator to evaluate its own factuality, which, as shown by recent works (Xu et al., 2024), suffers from self-bias and leads to inaccurate judgments.
@@
We further demonstrate how to leverage FenCE to improve the generator’s factuality with finetuning. To construct training data, as shown in Figure 1(b), we generate multiple responses for each prompt, use FenCE to judge and critique every claim in the responses, and replace false information with facts or remove it from the response, depending on whether the corresponding fact is within the generator’s knowledge (i.e., whether the generator outputs “unknown” when prompted about claim correctness). This mitigates introducing lesser-known facts into training data. Finally, unlike existing work that use the generator itself as the evaluator, we use FenCE to score each original and revised response and construct preference data, which reduces self-bias and produces more accurate judgments.
Experiments show that FenCE outperforms large open-source models such as Mistral-123B and strong proprietary models such as Claude-3 on LLM-Aggrefact. With FenCE, we improve Llama3-8B-chat’s factuality on FActScore (Min et al., 2023) by 14.45%, outperforming existing facutality finetuning methods by 6.96%. Analyses further show that after training with our recipe, the generator outputs less information for unfamiliar entities and outputs more information for popular ones, which suggests that it learns to only generate information that is likely to be factual.
Contributions. (1) We train a fine-grained critique-based evaluator, FenCE, by augmenting public datasets with textual critiques and more diverse source documents. (2) We propose a training recipe to improve the generator’s factuality by leveraging FenCE to improve and score its responses. (3) We conduct extensive experiments to validate both the judgment accuracy of FenCE and the factuality of the generator trained with FenCE, outperforming state-of-the-art factuality training methods by 6.96% on FActScore.
@@ 2.2 Improving Generator Factuality with FenCE
In this section, we use our evaluator, FenCE, to improve a generator model’s factuality, where we construct training data by revising and scoring the generator’s own responses. Compared to directly training the generator on factuality datasets, our method only requires a prompt set as inputs and hence enjors much better scalability.
@@
Proprietary models
[…]
@@
3 Evaluator Training Experiments
@@ 3.3 Result Analyses
4 Experiments on Improving Generator Factuality
We aim to answer the research questions: (RQ2) Can we leverage FenCE to improve the generator’s factuality? (RQ3) How well can our training recipe improve factuality compared to existing methods?
Model Name Test
@@ 4 Experiments on Improving Generator Factuality
Baselines (with Llama3-8B-chat as the evaluator)
@@
   SFT 20.05 18.13 52.52
   EVER 20.25 15.16 57.18
   FactTune 18.77 13.34 58.45
Ablations (with FenCE as the evaluator)
   SFT + FenCE 21.19 16.47 56.26
   Edit 20.68 14.42 58.91
   Coarse 20.07 12.89 60.89
   Edit + Coarse 20.03 11.09 64.37
   Edit/Remove + Coarse 20.40 10.79 65.41
Table 3: Results on the test split of FActScore. All the methods use Llama3-8B-chat as the base model. All the methods that revise the responses conduct at most 3 rounds of revision.
@@
4.1 Experimental Setup
Training and Evaluation Dataset. Following existing work (Kang et al., 2024a), we conduct experiments on the FActScore dataset (Min et al., 2023), which contain a set of celebrity names and the goal is to generate a biography with the prompt “Tell me a bio of <entity>.” We randomly split the unlabeled split of FActScore into 400 training prompts and 100 test prompts.
Baselines and Ablations. We implement three baselines: SFT, FactTune (Tian et al., 2024) and EVER (Kang et al., 2024a). All the methods sample $N$ candidate responses for each training prompt. SFT first uses the generator itself to score the responses by computing the percentage of factual claims and finetunes with the best response as target. FactTune first finetunes on all the candidates and then uses all $\binom{N}{2}$ candidate pairs as DPO pairs, preferring the one with higher score. EVER uses all $N$ candidates as rejected responses and constructs a preferred response by iteratively evaluating and correcting false information in each passage.
With respect to ablations, we first equip SFT, FactTune, and EVER with our evaluator, FenCE, denoted as “SFT + FenCE”, “Coarse”, and “Edit”, respectively. Then we implement “Edit + Coarse”, which corrects all false information without checking whether it corresponds to a lesser-known fact. We denote our full method as “Edit/Remove + Coarse”. More implementation details can be found in §A.
Evaluation Metric. We follow the FactScore procedure Min et al. (2023) to extract the number of correct and incorrect facts in each response, where we use Llama3-70B-chat to decompose the responses and use the same three tools in training to obtain source documents.
4.2 Main Results
Results in Table 3 show that our method significantly improves the percentage of facts of llama3-8B-chat by 14.45%, outperforming existing factuality finetuning methods by 6.96%. Furthermore, by using FenCE as the evaluator, we improve the performance of SFT, EVER, and FactTune by 1.73% $\sim$ 3.74%. This answers (RQ2) and shows FenCE’s ability to improve generators’ factuality.
By comparing with other ablations that use FenCE, we observe that our method of combining response editing/removing with coarse-level scoring achieves significantly better performance. Specifically, our method that only correct false information with common facts outperforms “Edit + Coarse”, which could introduce both common and lesser-known facts to the training data. The above results answer (RQ3), demonstrating the effectiveness of our training recipe.
@@ 4.3 Result Analyses
Similarly, we check the performance on claims describing different topics, where we first come up with a list of topics and then prompts Llama3-70B-chat to determine whether each claim covers any of the topics. As shown in Figure 5, our method also generates more factual claims under all the topics, and the performance gain is larger on unfamiliar topics (i.e., the topics where the generator produces lower percentage of facts).
Hyper-parameter Analysis. To investigate the effect of hyper-parameters, we first alternate the number of iterations to revise the responses and inspect the testing and training accuracy (i.e., the average % of facts of the best preferred responses). As shown in 6(a), with more revision iterations, we can always obtain preferred responses with better factuality, but the test performance converges after the third iteration. In other words, training data with fewer factual errors does not always transfer to better test performance.
In our experiments, we only choose the top-$k$ candidates (ranked by the percentage of facts) as preferred responses. We hence investigate the effect of $k$. As shown in 6(b), training on top-3 and top-5 responses leads to similar performance. Furthermore, with all the $k$ values in our experiments, our method consistently outperforms the “Edit + Coarse” ablation.
@@ 5 Related Work
Tool-Augmented LLMs. Tools have been shown to improve LLMs on diverse downstream tasks (Qin et al., 2023a; Schick et al., 2023; Qin et al., 2023b), including math reasoning (Parisi et al., 2022), question answering (Lazaridou et al., 2022), and multi-modal tasks (Tang et al., 2023). There have been works that apply tools to evaluation (Chern et al., 2023; Cho et al., 2023; Gou et al., 2024), including search engine, knowledge bases, etc. In this work, we train the evaluator on knowledge obtained by a multiplicity of tools to improve its ability to leverage tools in test time.
@@ 6 Conclusions
We improve LM generators’ factuality by training an open-source evaluator model, FenCE. To train FenCE to leverage diverse knowledge sources and to generate more informative feedback, we equip a combination of public datasets with textual critique along with judgment scores and obtain additional source documents by calling a multiplicity of tools. We then present a training recipe that leverages FenCE to finetune LM generators for better factuality, where we construct preference data by prompting FenCE to revise and score the generator’s responses, without introducing lesser-known facts in training. Experiments show that FenCE outperforms strong proprietary models such as Claude-3 on LLM-AggreFact. Our factuality training method improves Llama3-8B-chat’s factuality performance by 14.45% on FActScore.
@@ 8 Acknowledgement
We thank Xinran Zhao, Tim Franzmeyer, Zichun Yu, and Pengfei Liu for their helpful feedback on this work. This work was supported in part by NSF grant DSES 2222762.

### v3

@@ Abstract
Factuality evaluation aims to detect factual errors produced by language models (LMs) and hence guide the development of more factual models. Towards this goal, we train a factuality evaluator, FenCE, that provides LM generators with claim-level factuality feedback. In particular, we train FenCE to (1) generate textual critiques along with scores and (2) make claim-level judgment based on diverse source documents obtained by various tools, via data augmentation on a combination of public judgment datasets. We then present a framework that leverages FenCE to improve the factuality of LM generators by constructing training data. Specifically, we generate a set of candidate responses, ask FenCE to revise and score each response without introducing lesser-known facts, and train the generator by preferring highly scored revised responses. Experiments show that our data augmentation methods improve the evaluator’s accuracy by 2.9% on LLM-AggreFact. With FenCE, we improve Llama2-7B-chat/Llama3-8B-chat’s factuality rate by 16.86%/14.45% on FActScore, outperforming state-of-the-art factuality finetuning methods by 8.83%/6.96%.
@@ 1 Introduction
Alternatively, prior studies train the generator to output more factual responses, by preferring (1) generation candidates with higher factuality scores (Tian et al., 2024), which is limited by the generator’s capabilities, or (2) responses with false information corrected (Kang et al., 2024a), which is prone to introducing lesser-known facts. As shown in recent work (Gekhman et al., 2024; Ghosal et al., 2024), such preference training reinforces the model to generate information not well memorized during pretraining, and hence could even hurt factuality. Furthermore, the methods either leverage proprietary models that have restricted terms of use, or prompt the generator to evaluate its own factuality, which suffers from self-bias and leads to inaccurate judgments (Xu et al., 2024).
@@
We further demonstrate how to leverage FenCE to improve the generator’s factuality with finetuning. To construct training data (Figure 1(b)), we generate multiple responses for each prompt, use FenCE to judge and critique every claim in the responses, and replace false information with facts or remove it from the response, depending on whether the corresponding fact is within the generator’s knowledge (i.e., whether the generator outputs “unknown” when prompted about claim correctness). This mitigates introducing lesser-known facts into training data. Finally, unlike existing work that use the generator itself as the evaluator, we use FenCE to score each original and revised response and construct preference data, which reduces self-bias and produces more accurate judgments.
Experiments show that FenCE outperforms large open-source models such as Mistral-123B and strong proprietary models such as Claude-3 on LLM-Aggrefact. With FenCE, we improve Llama2-7B-chat’s factuality by 16.86% on FActScore (Min et al., 2023) and 17.64% on TruthfulQA Lin et al. (2022), outperforming existing factuality finetuning methods by 8.83% and 3.99%, respectively. Analyses further show that after training with our recipe, the generator outputs less information for unfamiliar entities and more information for popular ones, suggesting that it learns to only generate information that is likely to be factual.
Contributions. (1) We train a fine-grained critique-based evaluator, FenCE, by augmenting public datasets with textual critiques and more diverse source documents. (2) We propose a training recipe to improve the generator’s factuality by leveraging FenCE to improve and score its responses. (3) We conduct extensive experiments to validate both the judgment accuracy of FenCE and the factuality of the generator trained with FenCE, outperforming state-of-the-art factuality training methods by 8.83% on FActScore and 3.99% on TruthfulQA.
@@ 2.2 Improving Generator Factuality with FenCE
In this section, we use our evaluator, FenCE, to improve a generator model’s factuality, where we construct training data by revising and scoring the generator’s own responses. Compared to directly training the generator on factuality datasets, our method only requires a prompt set as inputs and hence enjoys much better scalability.
@@
Proprietary Models or Distilled Models
@@
3 Experiments on Evaluator Training
@@ 3.3 Result Analyses
4 Experiments on Generator Factuality
We aim to answer the research questions: (RQ2) Can we leverage FenCE to improve the generator’s factuality? (RQ3) How well can our training recipe improve the generator’s factuality?
4.1 Experimental Setup
Datasets and Evaluation Metric. Following prior works (Kang et al., 2024a; Zhang et al., 2024), we conduct experiments on FActScore (Min et al., 2023) and TruthfulQA (Lin et al., 2022). For FActScore, we randomly split the unlabeled split into 400 training and 100 test prompts. We compute the “Facts” metric by extracting correct and incorrect facts in each response, where we use Llama3-70B-chat to decompose the responses and use the same three tools in training to obtain source documents. For TruthfulQA, we randomly select 3 examples from each of the 38 categories as the training set and use the remaining 703 examples as the test set. We follow the “generation” setting and use the finetuned evaluator in the original paper to compute “True*Info”, the percentage of responses that are both truthful and informative.
Model Name FActScore TruthfulQA
# Facts # Errors % Facts % True*Info
Llama2-7B-chat 10.70 17.04 38.57 38.83
+ SFT 10.76 15.59 40.83 45.52
+ Self-Eval-SKT 11.02 14.18 43.73 48.65
+ EVER-Pref 11.24 15.11 42.66 51.07
+ FactTune-FS 11.23 12.87 46.60 52.48
[…]
+ E/R + Coarse 10.84 8.72 55.43 56.47
Llama3-8B-chat 17.83 17.16 50.96 58.89
+ SFT 20.05 18.13 52.52 59.17
+ Self-Eval-SKT 18.69 14.22 56.80 61.88
+ EVER-Pref 20.25 15.16 57.18 63.01
+ FactTune-FS 18.77 13.34 58.45 64.58
[…]
+ E/R + Coarse 20.40 10.79 65.41 67.14
Table 3: Comparison between our method and baselines on FActScore and TruthfulQA. “E/R” stands for “Edit/Remove”. All baselines use the zero-shot models as the evaluator in training and our method uses FenCE.
Model Name FActScore
@@ 4.1 Experimental Setup
(Ablations with FenCE as the evaluator)
@@
+ SFT + FenCE 21.19 16.47 56.26
+ Edit 20.68 14.42 58.91
+ Coarse 20.07 12.89 60.89
+ Edit + Coarse 20.03 11.09 64.37
[…]
+ E/R + Coarse 20.40 10.79 65.41
Table 4: Ablation study that compares different training recipes when equipped with FenCE as the evaluator. “SFT + FenCE”, “Edit”, and “Coarse” denote equipping “SFT”, “EVER-Pref”, and “FactTune-FS” in Table 3 with the FenCE evaluator.
@@
Baselines and Ablations. We implement four baselines: SFT, FactTune-FS (Tian et al., 2024), Self-Eval-SKT Zhang et al. (2024), and EVER-Pref (Kang et al., 2024a). All methods sample $N$ candidate responses for each training prompt. SFT first uses the generator itself to score responses by computing the percentage of factual claims and finetunes with the best response. FactTune-FS first finetunes on all candidates and then uses all $\binom{N}{2}$ candidate pairs as DPO pairs, preferring the one with a higher score based on retrieved context. Self-Eval-SKT self-trains an evaluator using the model’s own knowledge and uses the evaluator to score responses with no external context, also resulting in $\binom{N}{2}$ preference pairs. EVER-Pref uses all $N$ candidates as rejected responses and constructs a preferred response by iteratively evaluating and correcting false information in each passage.
For ablations, we first equip SFT, FactTune, and EVER with our evaluator, FenCE, and the three external tools, denoted as “SFT + FenCE”, “Coarse”, and “Edit” in Table 4, respectively. Then we implement “Edit + Coarse”, which corrects all false information without checking whether it corresponds to a lesser-known fact. We denote our full method as “E/R + Coarse” (Edit/Remove + Coarse). More implementation details can be found in §A.
Note that we do not compare to retrieval-augmented (RAG) methods Lewis et al. (2020) because it is not a fair comparison. RAG-based methods require access to external knowledge sources or tools in inference, while in our experiments, we only used “tell me a bio of <entity>” as the prompt. Furthermore, while most RAG-based methods aim to improve the understanding of the long retrieved context, our main focus is to train the model to better utilize the knowledge it has seen during pretraining.
We also do not compare to inference-time methods Li et al. (2024c); Chuang et al. (2024) because these methods need to call the model multiple times in inference, introducing heavy latency issues. In the real world, such latency is unacceptable in production. In comparison, our method uses the standard random decoding in inference and does not introduce extra inference-time cost. Furthermore, in principle, our method is orthogonal to those inference-time methods and can be combined.
@@
4.2 Main Results
Results in Table 3 answer (RQ2) and show that our method significantly improves Llama2/Llama3’s factuality by 16.85/14.45% on FActScore and 17.64%/8.25% on TruthfulQA. Results also show that our method significantly outperforms the best baseline on factuality training (e.g., by 8.83/6.96% on FActScore for Llama2/Llama3).
Table 4 presents the ablations with FenCE as the evaluator. The comparison between Table 3 and Table 4 shows that FenCE can improve the performance of SFT, EVER, and FactTune. For example, on FActScore, with Llama3 as the base model, SFT with FenCE as the evaluator outperforms SFT with Llama3 as the evaluator by 3.74%.
In Table 4, we observe that our method of combining response editing/removing with coarse-level scoring achieves significantly better performance. In particular, our full method, which only corrects false information with common facts, outperforms “Edit + Coarse”, which could introduce both common and lesser-known facts into the training data. The above results answer (RQ3), demonstrating the effectiveness of our training recipe.
@@ 4.3 Result Analyses
Similarly, we check the performance on claims describing different topics, where we first come up with a list of topics and then prompts Llama3-70B-chat to determine whether each claim covers any of the topics. In Figure 5, our method generates more factual claims under all the topics, and the performance gain is larger on unfamiliar topics (i.e., topics where the generator has lower scores).
Hyper-parameter Analysis. We first alternate the number of iterations to revise the responses and inspect the testing and training accuracy (i.e., the average % of facts of the best preferred responses). As shown in 6(a), with more revision iterations, we can always obtain preferred responses with better factuality, but the test performance converges after the third iteration. In other words, training data with fewer factual errors does not always transfer to better test performance.
In our experiments, we only choose the top-$k$ candidates (ranked by the percentage of facts) as preferred responses. We investigate the effect of $k$ in 6(b). We observe that training on top-3 and top-5 responses leads to similar performance. With all the $k$s in our experiments, our method consistently outperforms the “Edit + Coarse” ablation.
@@ 6 Conclusions
We improve LM generators’ factuality by training an open-source evaluator model, FenCE. To train FenCE to leverage diverse knowledge sources and to generate more informative feedback, we equip a combination of public datasets with textual critique along with judgment scores and obtain additional source documents by calling a multiplicity of tools. We then present a training recipe that leverages FenCE to finetune LM generators for better factuality, where we construct preference data by prompting FenCE to revise and score the generator’s responses, without introducing lesser-known facts in training. Experiments show that FenCE outperforms strong proprietary models such as Claude-3 on LLM-AggreFact. Our factuality training method improves Llama3-8B-chat’s factuality performance by 14.45% on FActScore and 8.25% on TruthfulQA.
@@ 8 Acknowledgement
We thank Xinran Zhao, Tim Franzmeyer, Zichun Yu, and Pengfei Liu for their helpful feedback on this work. This work was supported in part by NSF grant DSES-2222762. Yiqing Xie is supported by the Carnegie Mellon University Presidential Fellowship in the Language Technologies Institute.